OpenStudy (anonymous):

need some help with this one y/(y+2) - 1/(y-2) = 8/(y^2 - 4)

OpenStudy (dumbcow):

y = 5 combine fractions on left into single fraction then set numerators equal to each other solve for y

OpenStudy (anonymous):

i did that, didn't work

OpenStudy (dumbcow):

did you get y(y-2) -(y+2) = 8

OpenStudy (anonymous):

no, y-1/(y+2)(y-2) = 8/(y+2)(y-2)

OpenStudy (hunus):

\[y/(y+2) - 1/(y-2)-8/(y+2)(y-2)=0\] \[y/(y+2) - 1/(y-2)- 8/(y^2 - 4)=0\] \[y/(y+2) - 1/(y-2) -8/(y+2)(y-2)=0\] \[y(y-2)/(y+2)(y-2)-(y+2)/(y-2)(y+2)-8/(y-2)(y+2)=0\] \[(y^2-2y)/(y^2-4)-(y+2)/(y^2-4)-8/(y^2-4)=0\] \[(y^2-2y-y-2-8)/(y^2-4)=0\] \[(y^2-3y-10)/(y^2-4)=0\] \[((y-5)(y+2))/(y+2)(y-2)=0\] \[(y-5)/(y-2)=0\] \[y=5\]

OpenStudy (hunus):

Wrote out all of the steps so you could follow it more easily o.o

OpenStudy (dumbcow):

need common denominator to combine fractions multiply 1st fraction by (y-2)/(y-2) multiply 2nd fraction by (y+2)/(y+2) this way they have same denominator

OpenStudy (anonymous):

Hunus, I don't understand how you came up with y(y-2)/(y+2)(y-2) - (y-2)/(y-2)(y+2) - 8/(y-2)(y+2) so I'm stuck right there. Dumbcow, I already made the fractions have a common denominator. What about the other side of the equation though? Whatever is done to one side must be done to the other.

OpenStudy (hunus):

I made a common denominator by multiplying the fractions by \[(y-2)/(y-2) \] \[(y+2)/(y+2)\] respectively

OpenStudy (anonymous):

okay, now i'm stuck 2 steps down you have \[(y^{2}-2y-y-2-8)/(y^2-4)=0\] why isn't it\[(y ^{2}-2y-y+2-8)/(y^2-4)=0\]?

OpenStudy (hunus):

Because you're subtracting (y+2) -(y+2)=-y-2

OpenStudy (dumbcow):

because the neg sign must be distributed -(y+2) = -y -2

OpenStudy (anonymous):

okay

OpenStudy (hunus):

:D

OpenStudy (anonymous):

okay so you have (y-5)/(y-2)=0, how do you get, from that, that y=5?

OpenStudy (anonymous):

nothing can be done with that

OpenStudy (anonymous):

Multiply everything through by y-2.

OpenStudy (anonymous):

How will that work? The fractions won't have common denominators.

OpenStudy (anonymous):

what? \[\frac{y-5}{y-2} = 0\] so if you multiply by y-2, you get \[y-5 = 0(y-2) = 0\]

OpenStudy (anonymous):

:D okay, sorry about that, i was starting over thanks a lot